OpenStudy (anonymous):

If F(x) =2x/x-3 And FoG(x) =5x What is g(x)?

OpenStudy (faiqraees):

Fx x gx=5x Gx=5x/fx

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

that answer is kinda silly right?

OpenStudy (dumbcow):

\[f(g(x)) = \frac{2g(x)}{g(x) - 3} = 5x\] solve for g(x)

OpenStudy (dumbcow):

\[2g(x) = 5x(g(x) - 3)\] \[5x g(x) - 2 g(x) = 15x\] \[g(x) (5x-2) = 15x\] \[g(x) = \frac{15x}{5x-2}\]

OpenStudy (misty1212):

that should work i was going to suggest that \[F^{-1}\circ F(x)\circ G(x)=F^{-1}(5x)=G(x)\]

OpenStudy (misty1212):

\[F(x)=\frac{2x}{x-3}\\ F^{-1}(x)=\frac{3x}{x-2}\] \[F^{-1}(5x)=\frac{3\times 5x}{5x-2}\]